what is the slope of. 3x-2y=negative4? and how do you graph the linear equation?
Answer by ewatrrr(10682)

Hi
*Note: slope intercept form is y = mx + b where m is the slope and b the y-intercept.
.
3x -2y = -4
.
solve for y
y= (3/2)x + 2
.
slope = (3/2) (As y varies by 3, x varies in the same direction by 2!)
y-intercept is 2 (0,2) a point on the line
.
(0, 2) and (2,5) and (-2,-1) are points on the line
.